A body having S.H.M. of amplitude 5cm has a velocity of 50 cm/s when its displacement is 3cm .what is its period?

Grade:11

2 Answers

Sandeep Bhamoo
6 years ago
Vel. =dx/dt let ,x=5sin(wt). At x=3 sin(wt)=3/5 now velocity=w•cos(wt) =(2π/T)•4/5 -->. 50=8π/5T, T=8π/250. Please do approve my answer if correct
Amritpal Singh Saini
6 years ago
Dear student, first u can calculate the angular velocity from the equation v = w√(a^2-x^2) , from this u will get the w as 12.49 rad/s , also as for time period T=2 pi /w , on substitution of pi and w u will get the answer as 0.5028 sec . (Please do approve my answer)